a baby blue jay sits in a tall tree awaiting the arrival of its dinner. As the motherlands, on the nest, she drops a worm towards the hungry chicks mouth, but the worm misses and falls from the nest to the ground in 1.50 second. How high up is the nest?


1
Expert's answer
2021-09-23T08:30:38-0400
"h=\dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(1.50\ s)^2=11\ m."

Answer:

"h=11\ m."